In the Empirical Rule of data will fall in with two standard deviation.?
Approx 95% of the observations.
What the answer
If you spin a spinner 50 times how many times would you expect it to land on A?
The answer will depend on how man divisions the spinner has and how many of them are labelled A. Since you have chosen not to share this information the question cannot be answered.
If you spin a spinner 80 times how many times would you expect it to land on c?
The answer will depend on how man divisions the spinner has and how many of them are labelled c. Since you have chosen not to share this information the question cannot be answered.
If you flipped a coin 150 times about how many times would you expect it to land on heads?
A fair coin would be expected to land on heads 75 times.
What are odds of throwing a prime number with a dice?
The probability of eventually throwing a prime number is 1. On a single throw, of a fair die, the probability is 1/2.
If there is a correlation between Event A and Event B what can be a true statement?
Event B has something in common with Event A.
Who started the fidget spinner company?
"Catherine Hettinger has an inventor's mind: When she notices a problem, she tries to imagine a toy or device that can remedy it.
So when she heard about young boys throwing rocks at police officers and people walking past them while visiting her sister in Israel, her wheels started turning. She started brainstorming devices that could distract young children and provide them with a soothing toy to play with.
First, she thought of a soft rock that kids could throw. But then she tossed that idea aside, still thinking about other options when she returned to her home in Orlando, Fla. It was there that she eventually developed the idea that would become the original fidget spinner - more than two decades before the wildly popular device became the must-have toy for both kids and adults this year."
-h'ttp://time.com/money/4762207/fidget-spinner-inventor-catherine-hettinger
It is approx 0.41
If an event has a probability of 1 then it is?
Then it is a certainty that the event is 100% to happen.
What is the probability of a coin landing on either heads or tails?
Since it is a certainty that a coin must land on either heads or tails, the probability must be 1.
It is approx 28%.
0.55
What is the probability of drawing a club and a 9 from a deck of 52 playing cards?
The answer depends on whether or not the card is drawn randomly and also, whether or not the card drawn in the first attempt is replace.
Pr = ways_of_success/total_ways
Pr(Red) = 3/(5+3+1) = 3/9
Pr(Green) = 1/9
As they are independent:
Pr(Red or Green) = Pr(Red) + Pr(Green) = 3/9 + 1/9 = 4/9
----------------------------------------------------
Alternatively, as there are only three colours, selecting either Red or Green is the same as not selecting Blue:
Pr(Red or Green) = Pr(not Blue) = 1 - Pr(Blue)
→ Pr(Red or Green) = 1 - 5/9 = 4/9
Picking cards without replacing them are NOT independent events, since once you pick a card, you can't pick the same card again.
What is the probability of drawing a black 5 or a king of clubs from a standard deck of cards?
It is 5/52 for a single card, drawn randomly.
What is the probability of something ever happening?
Something just happened. As something would have to happen to make nothing happen, the probability of something else happening is approximately 100%.
If you have a more specific definition of something, probability can be calculated by [# of events] divided by [# of possible outcomes].
For example:
> The chances of rolling a 1 or a 2 on a 6-sided die
Is standard deviation biased or unbiased?
Standard deviation (SD) is neither biased nor unbiased. Estimates for SD can be biased but that depends on the formula used to calculate the estimate.
1. What I Would Do Is Add 2+3+5+6 Which Is 16
2. So Now You Know There Are 16 Marbles
3.Look At How Many Red And Blue Marbles There Are
4. Red=6
5.Blue=2
6. So 6+2=8
7. Your Answer Is 8/16 Or 1/2
8. So The Answer Is 8/16 Or In Simplest Form 1/2